Determin whether the graphs of the two lines are parallel.
y+3=9x
6x-y=-3
Are the graphs of the given equations parallel?
No or Yes

Answer by mananth(16946)

y+3=9x
6x-y=-3
Rewrite both equations in y= mx + b form and compare the co-efficients of x
y= 9x-3
y=6x+3
They are not same
so lines are not parallel